9. Infuse three 250 mL units of Packed Red Blood Cells (PRBCs) IV over 6 h. Use the following IV tubing label: List No. 14 Ror onty single do not use rostorize PLUMT BLOOD SET 100 mL Burette with Prepierced Injection Site. 170 Micron Filter Prepierced Port. Non-Vented 2 0 Answer gtt/min
Answer #1

Total volume of blood to be infused is 3× 250 = 750 ml

Total time = 6 hr = 6×60 = 360 mins

Drip factor of the blood transfusion set given =20 gtt /ml

Now, applying the formula :-

Drops /min = total volume to be infused / total time in mins× drop factor

= 750 / 360 × 20

= 41.6 ~42 drops / min = 42gtt /min